Find the constant of variation k if y varies directly with x, and y = 325 when x = -50.
Answer by Alan3354(69443)

Find the constant of variation k if y varies directly with x, and y = 325 when x = -50.
--------------------
y = kx
325 = k*(-50)
k = -325/50
k = -13/2